# PanoLAM: Large Avatar Model for Gaussian Full-Head Synthesis from One-shot Unposed Image

### v1

@@ 5 Conclusions
    Input    w/o coarse-to-fine  coarse output  fine output (Ours)
Figure 3: Comparison of our coarse-to-fine generation and direct dense generation. Without the coarse-to-fine strategy, the dense Gaussians are distributed on multiple layers, leading to degradation of rendering quality. Zoom in for details.
Figure 4: Comparison of different numbers of Gaussian points. A small number of Gaussian points cannot model texture details like hair strands and teeth. More high-fidelity Gaussian avatars are reconstructed with an increasing number of Gaussian points. Zoom in to see more details.
  Input   Renderings (2K Subj.)    Input   Renderings (165K Subj.)
Figure 5: Comparison of our network trained on different numbers of subjects. The model trained on a small number of subjects cannot generalize well to unseen identities.
Figure 6: Analysis of ray marching aggregation weights in the original Spherical Triplane in SphereHead. We cast rays and visualize the aggregation weights of each sampled feature in the original ray marching process. Two phenomena occur. Firstly, multiple peaks occur near the geometry surface hit by each ray in each curve, indicating that multiple layers of features should be aggregated rather than a single query to access the full valuable information residing in the spherical triplane. Secondly, each curve appears to be a different weight function. Thereafter, learning-based approaches, e.g., MLPs, can better model this aggregation operation.
Figure 7: Comparisons on different spherical triplane feature query and aggregation strategies. For each point, sampling only one sample of features from the spherical triplane cannot fully fetch all valuable features, resulting in a collapsed rendered image. Aggregating 32 points with Gaussian distribution weights gets better results, but still blurry. Our MLPs learn to aggregate features as the original NeRF and get quite better results.
Figure 8: Comparison of reconstruction and novel view synthesis of different methods on the testset.
Figure 9: Comparison of reconstruction and novel view synthesis of different methods on real-world in-the-wild images.
@@ Appendix A Appendix Overview
In this supplement, we present more visualization results in Sec. B, introduce our dataset generation and labeling process in Sec. C, discuss limitations of our work in Sec. D, and discuss ethical impacts in Sec. E.
Appendix B More Visualization Results
In this section, we show more results of our framework as well as more comparisons with previous work.
@@ Appendix B More Visualization Results
In Fig. 10, we render 360 images from our reconstructed Gaussian full-head given unposed images as input. As shown in the figure, after training on the proposed large-scale dataset, our framework can reconstruct Gaussian full-head and synthesize consistent novel view images.
@@
In Fig. 11, we show more results of different methods on the testset. Four views of images are rendered from the reconstructed 3D representation of different methods, showing the full-head quality. As is shown in the figure, our framework gets much higher fidelity results and has fewer artifacts in unseen regions compared to previous works.
@@
In Fig. 12, we show more results of different methods on real-world images sampled from the VFHQ dataset. As is shown in the figure, though trained on our synthesis only large-scale dataset, PanoLAM is able to generalize to the real-world images. Compared with previous methods, our framework reconstructs more texture details and maintains better identity fidelity.
Figure 10: Visualization of our Gaussian full-head synthesis from one-shot unposed image.
Figure 11: Comparison of reconstruction and novel view synthesis of different methods on the testset.
Figure 12: Comparison of reconstruction and novel view synthesis of different methods on in-the-wild real-world images.
Appendix C Generation and labeling process for our large-scale synthesis dataset.
@@
Given that EG3D is limited to rendering near-frontal images using its generated triplane NeRF, we sample images within a view range of approximately 72 from this model. Specifically, cameras are sampled from a pitch range of $\pm 26^{\circ}$ degrees and a yaw range of $\pm 36^{\circ}$ degrees relative to the front of the human face. To enhance the network’s adaptability to real-world images captured by various cameras, we vary the camera radius and focal length. The camera radius is drawn from a normal distribution centered at 2.7 with a standard deviation of 0.1, while the focal length is sampled from a normal distribution centered at 18.83 with a standard deviation of 1. Each triplane generates 32 random images. For SphereHead [43], we generate $360^{\circ}$ dataset. Specifically, we initially sample 8 views evenly spaced along the equatorial plane to ensure full-head coverage, dividing the 360-degree range. Additionally, we sample 24 images at random angles and apply the same randomization of camera radius and focal length as used in the EG3D sampling process. Visualization of some cases is shown in Fig. 13.
@@
Since the sampled data from 3D GANs may contain bad cases, which may hurt the training of the learning-based network. We also develop a labeling tool and remove bad cases manually to improve the dataset quality. Some examples of bad cases we removed are shown in Fig. 14.
Figure 13: Visualization of example cases in our sampled and cleaned datasets.
Figure 14: Examples of bad cases sampled from the 3D GANs prior, which are removed manually using our labeling tools.
Appendix D Limitations
Our dataset and model are built upon 3D GANs trained on different 2D datasets, e.g., FFHQ [34] for EG3D and the WildHead [43], CelebA [50], FFHQ, LPFF [79], and K-Hairstyle [38] datasets for SphereHead. The proposed dataset is sampled from these trained 3D GANs, and the network trained on it has similar biases to these datasets. The sampled datasets contain less Asian data and no cartoon heads, leading to worse reconstruction results on Asian faces and bad results on cartoon faces from the proposed framework. A possible solution to this problem is to train the 3D GANs on a more diverse 2D dataset for more diverse 3D head avatar sampling and 3D avatar reconstruction network training. Our model only handles static full-head reconstruction, and the generated head is not animatable. We leave this as future work.
Appendix E Ethical Impact

### v2

@@ Appendix A Appendix Overview
In this supplement, we present more results in Sec. B, more ablation studies in Sec. C, introduce our dataset generation and labeling process in Sec. E, discuss limitations of our work in Sec. D, and ethical statements in Sec. F. Sec. G illustrates the usage of LLM.
Appendix B More Results
In this section, we show more results as well as more comparisons with previous work.
Quantitative results on out-of-domain real-world datasets.
Table A1: Quantitative evaluation on real-world datasets.
[…]
PanoHead-PTI 24.37 0.806
SphereHead-PTI 24.65 0.811
LAM 18.72 0.713
Ours 27.45 0.882
We conduct experiments on in-the-wild real-world data to evaluate the generalizability of the proposed methods. We sample 9 identities from the VFHQ (Xie et al., 2022) test set and 9 identities from the Nersemble (Kirschstein et al., 2023) dataset as GaussianAvatar (Qian et al., 2024). We evaluate reconstruction quality using peak signal-to-noise ratio (PSNR) and structural similarity index (SSIM), following the protocol of GRAM‑HD (Xiang et al., 2023). Note that these data are not seen during training, which is suitable to evaluate the out-of-domain reconstruction capability of different methods. For each method, we synthesize 30 views of the same subject, train the surface reconstruction method NeuS2 (Wang et al., 2023b) on these images, and compute PSNR and SSIM on the resulting reconstructions. As shown in Table A1, PanoLam gets the best results on the reconstruction quality. While LAM (He et al., 2025) is trained on videos with a limited viewing angle, it cannot reconstruct the full head well, especially in the side and back views, and the results degrade largely in the $360^{\circ}$ evaluation. We also visualize some results in the right part of Fig. 4 and Fig. A4.
@@ Appendix B More Results
In Fig. A2, we render 360 images from our reconstructed Gaussian full-head given unposed images as input. As shown in the figure, after training on the proposed large-scale dataset, our framework can reconstruct Gaussian full-head and synthesize consistent novel view images.
@@
In Fig. A3, we show more results of different methods on the testset. Four views of images are rendered from the reconstructed 3D representation of different methods, showing the full-head quality. As shown in the figure, our framework gets much higher fidelity results and has fewer artifacts in unseen regions compared to previous works.
@@
In Fig. A4, we show more results of different methods on real-world images sampled from the VFHQ dataset. As shown in the figure, though trained on our synthesis only large-scale dataset, PanoLAM is able to generalize to the real-world images. Compared with previous methods, our framework reconstructs more texture details and maintains better identity fidelity.
Appendix C More Ablation
Effect of Different Number of Gaussian Points.
Figure A1: Comparison of different numbers of Gaussian points. Zoom in to see more details.
We ablate the effect of different numbers of Gaussian points with our coarse-to-fine strategy in Fig. A1. As shown in the figure, a small number of Gaussian points cannot model texture details like hair strands and teeth, leading to blurry results in these regions. In contrast, more high-fidelity Gaussian avatars are reconstructed with an increasing number of Gaussian points. However, directly increasing the number of Gaussian points causes the problem of insufficient optimization of Gaussians and out-of-memory issues. This further validate the effectiveness of our proposed coarse-to-fine training strategy.
Appendix D Limitations
Our dataset and model are built upon 3D GANs trained on different 2D datasets, e.g., FFHQ (Kazemi & Sullivan, 2014) for EG3D and the WildHead (Li et al., 2024), CelebA (Liu et al., 2015), FFHQ, LPFF (Wu et al., 2023), and K-Hairstyle (Kim et al., 2021) datasets for SphereHead. The proposed dataset is sampled from these trained 3D GANs, and the network trained on it has similar biases to these datasets. The sampled datasets contain less Asian data and no cartoon heads, leading to worse reconstruction results on Asian faces and bad results on cartoon faces from the proposed framework. A possible solution to this problem is to train the 3D GANs on a more diverse 2D dataset for more diverse 3D head avatar sampling and 3D avatar reconstruction network training. Our model only handles static full-head reconstruction, and the generated head is not animatable. We leave this as future work.
Appendix E Generation and labeling process for our large-scale synthesis dataset.
@@ Appendix D Limitations
Given that EG3D is limited to rendering near-frontal images using its generated triplane NeRF, we sample images within a view range of approximately 72 from this model. Specifically, cameras are sampled from a pitch range of $\pm 26^{\circ}$ degrees and a yaw range of $\pm 36^{\circ}$ degrees relative to the front of the human face. To enhance the network’s adaptability to real-world images captured by various cameras, we vary the camera radius and focal length. The camera radius is drawn from a normal distribution centered at 2.7 with a standard deviation of 0.1, while the focal length is sampled from a normal distribution centered at 18.83 with a standard deviation of 1. Each triplane generates 32 random images. For SphereHead (Li et al., 2024), we generate $360^{\circ}$ dataset. Specifically, we initially sample 8 views evenly spaced along the equatorial plane to ensure full-head coverage, dividing the 360-degree range. Additionally, we sample 24 images at random angles and apply the same randomization of camera radius and focal length as used in the EG3D sampling process. Visualization of some cases is shown in Fig. A5.
@@
Since the sampled data from 3D GANs may contain bad cases, this may hurt the training of the learning-based network. We also develop a labeling tool and remove bad cases manually to improve the dataset quality. Some examples of bad cases we removed are shown in Fig. A6.
Appendix F Ethical Impact
@@
Appendix G State of LLM Usage
We use LLM to assist with paper polishing on grammar. Each sentence polished by LLM is checked to express our original meaning. There is no further use of LLM for the idea formulation, experiments, coding, and main paper writing.
Figure A2: Visualization of our Gaussian full-head synthesis from one-shot unposed image.
Figure A3: Comparison of reconstruction and novel view synthesis of different methods on the testset.
Figure A4: Comparison of reconstruction and novel view synthesis of different methods on in-the-wild real-world images.
Figure A5: Visualization of example cases in our sampled and cleaned datasets.
Figure A6: Examples of bad cases sampled from the 3D GANs prior, which are removed manually using our labeling tools.